A cube of side 4 cm is cut into cubes of side 1 cm, then total surface area of all the small cubes is ______ .

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Answer: Explaination: 384 cm² Volume of bigger cube = 4 × 4 × 4 = 64 cm3 Volume of one smaller cube = 1 × 1 × 1 = 1 cm3 Number of smaller cubes = 641 = 64 Surface area of one smaller cube = 6 × 1² = 6 cm² ∴ Total surface area = 6 × 64 = 384 cm²